One day the teperature at 8am was 41*F, and by 1pm the temperature was 61*F. What was the hourly rate of temperature change
Answer by palanisamy(496)

One day the teperature at 8am was 41*F, and by 1pm the temperature was 61*F
In 5 hours the change in temperature = 61-41 = 20
In 1 hour the change in temperature = 20/5 = 4
Therefore hourly rate of temperature change = 4*F